# Soot Planets instead of Water Worlds

### v1

@@ 4 Implications
Some characteristics of the interiors of soot-rich planets may inhibit habitability. Abundant diamond in a solidified silicate mantle could increase interior viscosity and thermal conductivity by several orders of magnitude, thus retarding mantle convection and volatile cycling, and potentially yielding inhospitable surface conditions (C. T. Unterborn et al., 2014). A carbon-rich core may not provide the buoyancy needed to support a global magnetic field, which protects life from stellar wind and cosmic radiation, though other dynamo mechanisms are possible in a sub-Neptune that retains a partially molten deep silicate interior (L. Stixrude et al., 2021).
Alternately, soot-rich planets can have favorable consequences for habitability and the development of life. Persistent supplies of methane and other reduced gases, including likely significant partial pressures of NH3 and H2 P. Liggins et al. (2022), are anticipated as volcanogenic products from the thermally processed interiors of soot-rich planets (E. A. Bergin et al., 2023). Abundant methane with reducing atmospheric and oceanic conditions are possibly a prerequisite for progression of prebiotic chemistry (T. M. McCollom, 2013). An instructive analog may be found on Titan, where hydrolysis of methane-produced hazes have been shown via laboratory experiments to form simple amino acids (H. J. Cleaves et al., 2014). Although reactions between organics and liquid water may not occur on Titan (C. Neish et al., 2024), organics could be created in the atmosphere of soot-water worlds and provide a consistent source to potentially foster a biosphere in underlying oceans.
@@
Acknowledgment The research shown here acknowledges use of the Hypatia Catalog Database, an online compilation of stellar abundance data as described in N. R. Hinkel et al. (2014), which was supported by NASA’s Nexus for Exoplanet System Science (NExSS) research coordination network and the Vanderbilt Initiative in Data-Intensive Astrophysics (VIDA). JL acknowledges support from National Science Foundation (NSF) grant EAR2317024. MMH acknowledges support from NSF grant EAR2246915 and National Aeronautics and Space Administration (NASA) grant 80NSSC19K0959. EAB and GAB acknowledge support from NASA grant 80NSSC24K0149. EM-RK acknowledges funding from AEThER via the Alfred P. Sloan Foundation under grant G202114194. FJC and EAB acknowledge support from NASA Award 80NSSC20K0259. This project is supported, in part, by funding from Two Sigma Investments, LP to EAB. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflects the views of Two Sigma Investments, LP.
Code Availability Statement The code used to calculate M-R relations is made available on Zenodo by Li, J. (2025). Soot Planet Code:Zenodo: https://doi.org/10.5281/zenodo.16101631
@@ Appendix A Model Planet Compositions
Guided by the solar, meteoritic, and cometary compositions, together with first-order constraints from planet formation models (J. Li et al., 2021), we constructed model planets from four chemical components: rock, soot, water, and hydrogen (Table A1). The rock component contains four elements, oxygen (O), magnesium (Mg), silicon (Si), and iron (Fe), with the solar Mg/Si and Fe/Si ratios (K. Lodders, 2010). It consists of two phases, with all Mg and Si combined with O to form MgSiO3, and all Fe present as metal. The soot component is assumed to have C:H:O atomic ratio of 100:77±2:14±3, the average value of insoluble organic matter (IOM) in primitive chondrites and comet Halley (C. Alexander et al., 2017). It consists of multiple phases, including meteoritic insoluble organic matter (IOM) and carbonaceous materials found in comets. The water component and hydrogen component are taken to be pure H2O and pure H2, respectively. The specific carriers of the carbon are a variety of materials, including species identified above, which we term “refractory organics.”
@@
The C/H ratios vary from 0.1-0.2 in Murchison meteorites, 0.3-0.5 in Bennu, and 0.4 in Tagish Lake (D. P. Glavin et al., 2025), to 1.7 in Ryugu(T. Yokoyama et al., 2025). We adopted the C/H ratio of 1.3 for IOM (C. Alexander et al., 2017). This ratio affects the relative amounts of soot and water in the models of soot planets and soot-water worlds, but does not change our conclusions.
@@
We consider three types of model planets with four distinct chemical compositions: a rocky planet, a soot planet, a dry version of soot-water world, and a wet version of soot-water world(Table A1). The model rocky planet only contains the rock component. The model soot planet contains the rock and soot components and has a C/Si ratio at half of the solar value (K. Lodders, 2010), following the assumption that half of the solar carbon is preserved as refractory organics and the other half is lost as carbon monoxide (CO) (J. Li et al., 2021). This assumption is consistent with the observation that on average dust is half rock and half soot(M. Rubin et al., 2019). Hirschman et al. (M. M. Hirschmann et al., 2021) showed that devolatilization during small-body differentiation is a key process in shaping the volatile inventory of terrestrial planets derived from planetesimals and planetary embryos. We thus consider a range of dust-to-ice ratio to account for variable degrees of volatile loss by parent-body processing during the formative stages of planets. For the soot-water worlds, we consider 1 as a lower bound on the dust-to-ice ratio and 3 as an upper bound, as observed in Comet 67P (M. Rubin et al., 2019; U. Malamud et al., 2024)(U. Malamud & D. Prialnik, 2015; C. J. Bierson & F. Nimmo, 2019). Accordingly, the model soot planet contains about one quarter of soot and three quarters of rock by mass, and the model dry and wet soot-water worlds consist of about $\frac{1}{2}$rock-$\frac{1}{5}$soot-$\frac{1}{4}$water and $\frac{1}{3}$rock-$\frac{1}{7}$soot-$\frac{1}{2}$water, respectively (Table A1).
@@
The compositions of real planets likely deviate from these models. For instance, inclusion of hydrous minerals may alter a planet’s H/C ratio, as H2O can be stored in hydrous minerals such as clays and serpentine in chondritic meteorites and asteroids (e.g.(B. Bitsch et al., 2019a; V. E. Hamilton et al., 2019)). Formation of hydrous phases requires reaction between anhydrous precursors and water ice and therefore hydrous phases are abundant only beyond water snow line. Once formed, hydrous phases may be delivered to the inner disks through radial migration and they may survive the high temperatures due to sluggish dehydration (F. Ciesla & D. Lauretta, 2005). Incorporating hydrous minerals would increase the H/C ratios of rocky or soot planets. On the other hand, hydrous phases are thought to form via parent body alteration, during which water loss is expected, while soot is expected to survive the process, which occurs below the decomposition temperatures of the soot. As a result, the presence of hydrous minerals in soot-water worlds may be associated with lower H/C ratios. Finally, the required parent body processing also implies that a planet accreting from pebbles likely contains little hydrous materials.
@@
1 With Si fixed at 1 K. Lodders (2003) “–” denotes negligible.
@@
Note that the previous water worlds (R. Luque & E. Pallé, 2022) consist of rock and water, whereas our model soot-water worlds contain soot as a major component, as it is unlikely that a planet can accrete appreciable water without also accreting soot (J. Li et al., 2021). Carbon-rich compounds are known to significantly contribute to the mass budget of a variety of objects in the Solar System including comets, carbonaceous chondrites, icy moons, TNOs, and ice giants, so they should contribute to the composition of planets around other stars, so long as they coalesced outside the “soot line”. Combining chemical data from materials in the solar system (e.g., comets and chondritic meteorites) and thermochemical modeling, our study demonstrates that water worlds are also expected to contain soot, and that planets rich in “soot” (refractory organic carbon) can account for the observed exoplanets with low average densities that were previously attributed to a binary combination of rock and water.
Because soot consists largely of hydrocarbons, the interiors of planets rich in soot will also be comparatively reduced and produce secondary atmospheres rich in methane (E. A. Bergin et al., 2023). Reactions such as 6 CH (in soot) + Fe2O3 = C + 2FeO + 3H2O will effectively neutralize any oxidized iron so long as the soot/oxidized iron mass ratio is greater than about 0.3. In soot planets the total soot mass is more than 20% and the majority of iron is accreted as metal, and therefore volcanogenic C is outgassed as a combination of CH4 and CO.
@@ Appendix B C/O Ratio
There is a growing literature on the overall carbon to oxygen content of exoplanetary atmospheres along side their potential link to the disk composition, all in the context of an assumed/known stellar C/O ratio (K. I. Öberg et al., 2011b; N. Madhusudhan, 2012; T. P. Greene et al., 2016; J. I. Moses et al., 2013; N. Madhusudhan, 2019; E. A. Bergin et al., 2024). The overall context is described by (K. I. Öberg et al., 2011b) where the stellar C/O is assumed to be solar and the disk composition reflects that the major carriers of elemental C and O based on ISM constraints. These are silicate dust, H2O, CO, and CO2 for oxygen and carbonaceous dust, CO, and CO2 for carbon. This particular model mostly focused on the volatile ice components that condense $\geq$1 au from the star (i.e. CO, CO2, and H2O). This has relevance for the composition of gas giants in the context of the core-accretion model of giant planet formation as discussed in the references above.
@@
The model discussed in this paper is an extension of this concept, but with a focus on the innermost region of protoplanetary disks from the water ice line and inwards. This includes the sublimation fronts of soot (carbonaceous dust) (M. E. Kress et al., 2010; J. Li et al., 2021) and silicates. It is this region that matters most for rocky planets (or the cores of mini-Neptune’s) as discussed by J. Li et al. (2021) and E. A. Bergin et al. (2023). Additional motivational work is found in Bond et al. (J. C. Bond et al., 2011) and Moriarty et al.(J. Moriarty et al., 2014) who use thermochemical equilibrium simulations to argue that carbon-rich bodies could be present in the extra-solar inventory (see also C. J. Shakespeare et al., 2025). However, the formation of macromolecular hydrocarbons is not a product of equilibrium condensation (K. Lodders, 2003). Thus, the concept of the irreversible destruction of soot via sublimation (J. Li et al., 2021) places a new constraint on the location of carbon-rich material with attendant implications for planet formation.
@@
In this regard, a central question is how common is the C/O ratio of the solar system in comparison to exoplanetary systems. N. Madhusudhan et al. (2012) explored the possibility of a carbon-rich interior for the super-Earth 55 Cancri under the assumption of a stellar C/O >> 1. This composition directly implies the presence of excess carbon in the disk. If this excess carbon could be placed in refractory solids, then this would potentially be reflected in the composition of the rocky planet (A. Johansen & M. Lambrechts, 2017). In Fig. A1 we present a compilation of stellar C/O ratios for stars that host known planetary systems with massive planets (M $\geq$30 M; shown in Red) and low mass planets (shown in Blue). As can be seen here, for both cases, the vast majority of systems have near solar C/O ratios. This is consistent with other analysis (P. E. Nissen, 2013; R. da Silva et al., 2024) and the argument that systems with extreme C/O ratios (>> 1) are rare in the nearby stellar population (J. J. Fortney, 2012).
@@
Figure A1: Histogram of stellar C/O ratios, (C/O), for planet host stars as compiled by the Hypatia Catalog (N. R. Hinkel et al., 2014). The histogram shown in red (High Mass Planets) isolates systems where the most massive planet is $\geq$30 M and the one is blue (Low Mass Planets) isolates those with all planet mass below that value. The median error for each sample is shown with its magnitude comparable to the length of the line. The solar value is also show as reference using oxygen abundance estimates by Bergemann et al. (M. Bergemann et al., 2021) and oxygen by Asplund et al. (M. Asplund et al., 2021).
@@ Appendix C Soot Density
To estimate the density of the soot component at ambient pressure and temperature conditions (100 kPa = 1 bar, 300 K), we compiled the densities of an array of planet-forming materials, including iron metal, silicates, water ice, and various carbon-bearing phases including diamond, graphite, carbonates, and iron carbides (Table A2). The compilation also includes a variety of ices observed at cryogenic conditions as well metastable phases which form at the high pressures prevailing in planetary interiors. Due to thermal expansion upon heating from 80 K to 195 K, the densities of H2O1h ice and CO2I ice decrease by 0.7% and 8%, respectively (T. P. Mangan et al., 2017). At 300 K, the densities of fictitious H2O1h ice and CO2I ice are 0.896 g/cc and 1.398 g/cc, respectively, after correcting for thermal expansion.
@@
The densities of the selected phases at 100 kPa and 300 K show a linear correlation with their mean atomic numbers ($\overline{Z}$) (Fig. 2). As expected, high-pressure polymorphs such as water ice VII, diamond, bridgmanite, and Fe7C3 iron carbide plot above the linear trend. Most cryogenic ices and simple organic materials follow the linear correlation, with a few outliers falling below the trend (Fig. 2 inset). A least-squares fit of eleven stable phases at ambient conditions (Table A2) yields $\rho_{0}$ = 0.327 $\overline{Z}$, with R2 = 0.99, or $\rho_{0}$ = 0.316 $\overline{Z}$ + 0.161, with R2 = 0.95. The $\overline{Z}$ of soot with C:H:O = 100:78:18 in atomic ratio is 4.17 and has a estimated density of 1.32(6) g/cc at 100 kPa and 300 K based on the adopted regression of $\rho_{0}$ = 0.317(15) $\overline{Z}$. This value is comparable to the proposed density of carbonaceous matter (1.4 g/cc) that coexists with hydrous silicates at 100 kPa and up to 900 K (A. Néri et al., 2020).

### v2

@@ 4 Implications
Some characteristics of the interiors of soot-rich planets may inhibit habitability. Abundant diamond in a solidified silicate mantle could increase interior viscosity and thermal conductivity by several orders of magnitude, thus retarding mantle convection and volatile cycling, and potentially yielding inhospitable surface conditions (Unterborn2014). A carbon-rich core may not provide the buoyancy needed to support a global magnetic field, which protects life from stellar wind and cosmic radiation, though other dynamo mechanisms are possible in a sub-Neptune that retains a partially molten deep silicate interior (Stixrude2021).
Alternately, soot-rich planets can have favorable consequences for habitability and the development of life. Persistent supplies of methane and other reduced gases, including likely significant partial pressures of NH3 and H2 liggins2022, are anticipated as volcanogenic products from the thermally processed interiors of soot-rich planets (Bergin2023). Abundant methane with reducing atmospheric and oceanic conditions are possibly a prerequisite for progression of prebiotic chemistry (McCollom13). An instructive analog may be found on Titan, where hydrolysis of methane-produced hazes have been shown via laboratory experiments to form simple amino acids (Cleaves2014). Although reactions between organics and liquid water may not occur on Titan (Neish2024), organics could be created in the atmosphere of soot-water worlds and provide a consistent source to potentially foster a biosphere in underlying oceans.
@@
Acknowledgment The research shown here acknowledges use of the Hypatia Catalog Database, an online compilation of stellar abundance data as described in Hypatia, which was supported by NASA’s Nexus for Exoplanet System Science (NExSS) research coordination network and the Vanderbilt Initiative in Data-Intensive Astrophysics (VIDA). JL acknowledges support from National Science Foundation (NSF) grant EAR2317024. MMH acknowledges support from NSF grant EAR2246915 and National Aeronautics and Space Administration (NASA) grant 80NSSC19K0959. EAB and GAB acknowledge support from NASA grant 80NSSC24K0149. EM-RK acknowledges funding from AEThER via the Alfred P. Sloan Foundation under grant G202114194. FJC and EAB acknowledge support from NASA Award 80NSSC20K0259. This project is supported, in part, by funding from Two Sigma Investments, LP to EAB. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflects the views of Two Sigma Investments, LP.
Code Availability Statement The codes for calculating the M-R relations are made available on Zenodo (li25Zenodo).
This study investigates the potential role of carbon-rich compounds in the interiors of exoplanets. There are considerable uncertainties in the internal structure and the thermal profiles of the modeled planets as well as in the distribution and forms of the soot component in exoplanets. We derive bounds on the mass-radius relations by treating the soot component as a fictive phase contributing to the properties of the planets, and then estimating the density of the soot component on the basis of a broad correlation between density and mean atomic number. We found that the uncertainties in the forms of carbon in exoplanets and the equations-of state (EoS) of relevant phases are comparable to the observational uncertainties in mass and radius; therefore, our simplifying approach allows meaningful estimates to examine whether soot planets can account for observed mass-radius relations.
@@ Appendix A Model Planet Compositions
Guided by the solar, meteoritic, and cometary compositions, together with first-order constraints from planet formation models (Li21), we constructed model planets from four chemical components: rock, soot, water, and hydrogen (Table A1). The rock component contains four elements, oxygen (O), magnesium (Mg), silicon (Si), and iron (Fe), with the solar Mg/Si and Fe/Si ratios (Lodders10). It consists of two phases, with all Mg and Si combined with O to form MgSiO3, and all Fe present as metal. The soot component is assumed to have C:H:O atomic ratio of 100:77±2:14±3, the average value of insoluble organic matter (IOM) in primitive chondrites and comet Halley (Alexander17). It consists of multiple phases, including meteoritic insoluble organic matter (IOM) and carbonaceous materials found in comets. The water component and hydrogen component are taken to be pure H2O and pure H2, respectively. The specific carriers of the carbon are a variety of materials, including species identified above, which we term “refractory organics.”
@@
The assumed C/Si ratio is consistent with the observation that on average dust is half rock and half soot (Rubin19) and is also consistent with C/Si ratios detected in some polluted white dwarfs (koester2014atmospheric; 2024A&A...691A.352W), in comets 67P (Bardyn2017), Halley (Fomenkova1999), and other comets (Woodward2021). The C/H ratios vary from 0.1-0.2 in Murchison meteorites, 0.3-0.5 in Bennu, and 0.4 in Tagish Lake (2025NatAs...9..199G), to 1.7 in Ryugu(2025GeocJ..59...45Y). We adopted the C/H ratio of 1.3 for IOM (Alexander17). This ratio affects the relative amounts of soot and water in the models of soot planets and soot-water worlds, but does not change our conclusions.
@@
We consider three types of model planets with four distinct chemical compositions: a rocky planet, a soot planet, a dry version of soot-water world, and a wet version of soot-water world(Table A1). The model rocky planet only contains the rock component. The model soot planet contains the rock and soot components and has a C/Si ratio at half of the solar value (Lodders10), following the assumption that half of the solar carbon is preserved as refractory organics and the other half is lost as carbon monoxide (CO) (Li21). This assumption is consistent with the observation that on average dust is half rock and half soot(Rubin19) and is also consistent with C/Si ratios detected in some polluted white dwarfs (koester2014atmospheric; 2024A&A...691A.352W), in comets 67P (Bardyn2017), Halley (Fomenkova1999), and numerous others (Woodward2021). Hirschmann et al. (Hirschmann2021PNAS..11826779H) showed that devolatilization during small-body differentiation is a key process in shaping the volatile inventory of terrestrial planets derived from planetesimals and planetary embryos. We thus consider a range of dust-to-ice ratio to account for variable degrees of volatile loss by parent-body processing during the formative stages of planets. For the soot-water worlds, we consider 1 as a lower bound on the dust-to-ice ratio and 3 as an upper bound, as observed in Comet 67P (Rubin19; Malamud2024)(2015Icar..246...21M; 2019Icar..326...10B). Accordingly, the model soot planet contains about one quarter of soot and three quarters of rock by mass, and the model dry and wet soot-water worlds consist of about $\frac{1}{2}$rock-$\frac{1}{5}$soot-$\frac{1}{4}$water and $\frac{1}{3}$rock-$\frac{1}{7}$soot-$\frac{1}{2}$water, respectively (Table A1).
@@
The compositions of real planets likely deviate from these models. For instance, inclusion of hydrous minerals may alter a planet’s H/C ratio, as H2O can be stored in hydrous minerals such as clays and serpentine in chondritic meteorites and asteroids (e.g.(Bitsch2019; 2019NatAs...3..332H)). Formation of hydrous phases requires reaction between anhydrous precursors and water ice and therefore hydrous phases are abundant only beyond water snow line. Once formed, hydrous phases may be delivered to the inner disks through radial migration and they may survive the high temperatures due to sluggish dehydration (Ciesla2005E&PSL.231....1C). Incorporating hydrous minerals would increase the H/C ratios of rocky or soot planets. On the other hand, hydrous phases are thought to form via parent body alteration, during which water loss is expected, while soot is expected to survive the process, which occurs below the decomposition temperatures of the soot. As a result, the presence of hydrous minerals in soot-water worlds may be associated with lower H/C ratios. Finally, the required parent body processing also implies that a planet accreting from pebbles likely contains little hydrous materials.
@@
1 With Si fixed at 1 Lodders03 “–” denotes negligible.
@@
Note that the previous water worlds (Luque22) consist of rock and water, whereas our model soot-water worlds contain soot as a major component, as it is unlikely that a planet can accrete appreciable water without also accreting soot (Li21). Carbon-rich compounds are known to significantly contribute to the mass budget of a variety of objects in the Solar System including comets, carbonaceous chondrites, icy moons, TNOs, and ice giants, so they should contribute to the composition of planets around other stars, so long as they coalesced outside the “soot line”. Combining chemical data from materials in the solar system (e.g., comets and chondritic meteorites) and thermochemical modeling, our study demonstrates that water worlds are also expected to contain soot, and that planets rich in “soot” (refractory organic carbon) can account for the observed exoplanets with low average densities that were previously attributed to a binary combination of rock and water.
Because soot consists largely of hydrocarbons, the interiors of planets rich in soot will also be comparatively reduced and produce secondary atmospheres rich in methane (Bergin2023). Reactions such as 6 CH (in soot) + Fe2O3 = C + 2FeO + 3H2O will effectively neutralize any oxidized iron so long as the soot/oxidized iron mass ratio is greater than about 0.3. In soot planets the total soot mass is more than 20% and the majority of iron is accreted as metal, and therefore volcanogenic C is outgassed as a combination of CH4 and CO.
@@ Appendix B C/O Ratio
There is a growing literature on the overall carbon to oxygen content of exoplanetary atmospheres along side their potential link to the disk composition, all in the context of an assumed/known stellar C/O ratio (Oberg11_C_O; 2012ApJ...758...36M; 2016ApJ...817...17G; 2013ApJ...763...25M; 2019ARA&A..57..617M; 2024ApJ...969L..21B). The overall context is described by (Oberg11_C_O) where the stellar C/O is assumed to be solar and the disk composition reflects that the major carriers of elemental C and O based on ISM constraints. These are silicate dust, H2O, CO, and CO2 for oxygen and carbonaceous dust, CO, and CO2 for carbon. This particular model mostly focused on the volatile ice components that condense $\geq$1 au from the star (i.e. CO, CO2, and H2O). This has relevance for the composition of gas giants in the context of the core-accretion model of giant planet formation as discussed in the references above.
@@
The model discussed in this paper is an extension of this concept, but with a focus on the innermost region of protoplanetary disks from the water ice line and inwards. This includes the sublimation fronts of soot (carbonaceous dust) (Kress2010; Li21) and silicates. It is this region that matters most for rocky planets (or the cores of mini-Neptune’s) as discussed by Li21 and Bergin2023. Additional motivational work is found in Bond et al. (Bond11) and Moriarty et al.(2014ApJ...787...81M) who use thermochemical equilibrium simulations to argue that carbon-rich bodies could be present in the extra-solar inventory (see also 2025AJ....169..180S). However, the formation of macromolecular hydrocarbons is not a product of equilibrium condensation (Lodders03). Thus, the concept of the irreversible destruction of soot via sublimation (Li21) places a new constraint on the location of carbon-rich material with attendant implications for planet formation.
@@
In this regard, a central question is how common is the C/O ratio of the solar system in comparison to exoplanetary systems. Madhu12 explored the possibility of a carbon-rich interior for the super-Earth 55 Cancri under the assumption of a stellar C/O >> 1. This composition directly implies the presence of excess carbon in the disk. If this excess carbon could be placed in refractory solids, then this would potentially be reflected in the composition of the rocky planet (Johansen2017). In Fig. A1 we present a compilation of stellar C/O ratios for stars that host known planetary systems with massive planets (M $\geq$30 M; shown in Red) and low mass planets (shown in Blue). As can be seen here, for both cases, the vast majority of systems have near solar C/O ratios. This is consistent with other analysis (Nissen2013; 2024A&A...688A.193D) and the argument that systems with extreme C/O ratios (>> 1) are rare in the nearby stellar population (2012ApJ...747L..27F).
@@
Figure A1: Histogram of stellar C/O ratios, (C/O), for planet host stars as compiled by the Hypatia Catalog (Hypatia). The histogram shown in red (High Mass Planets) isolates systems where the most massive planet is $\geq$30 M and the one is blue (Low Mass Planets) isolates those with all planet mass below that value. The median error for each sample is shown with its magnitude comparable to the length of the line. The solar value is also show as reference using oxygen abundance estimates by Bergemann et al. (2021MNRAS.508.2236B) and oxygen by Asplund et al. (2021A&A...653A.141A).
@@ Appendix C Soot Density
To estimate the density of the soot component at ambient pressure and temperature conditions (100 kPa = 1 bar, 300 K), we compiled the densities of an array of planet-forming materials, including iron metal, silicates, water ice, and various carbon-bearing phases including diamond, graphite, carbonates, and iron carbides (Table A2). The compilation also includes a variety of ices observed at cryogenic conditions as well metastable phases which form at the high pressures prevailing in planetary interiors. Due to thermal expansion upon heating from 80 K to 195 K, the densities of H2O1h ice and CO2I ice decrease by 0.7% and 8%, respectively (Mangan2017). At 300 K, the densities of fictitious H2O1h ice and CO2I ice are 0.896 g/cc and 1.398 g/cc, respectively, after correcting for thermal expansion.
@@
The densities of the selected phases at 100 kPa and 300 K show a linear correlation with their mean atomic numbers ($\overline{Z}$) (Fig. 2). As expected, high-pressure polymorphs such as water ice VII, diamond, bridgmanite, and Fe7C3 iron carbide plot above the linear trend. Most cryogenic ices and simple organic materials follow the linear correlation, with a few outliers falling below the trend (Fig. 2 inset). A least-squares fit of eleven stable phases at ambient conditions (Table A2) yields $\rho_{0}$ = 0.327 $\overline{Z}$, with R2 = 0.99, or $\rho_{0}$ = 0.316 $\overline{Z}$ + 0.161, with R2 = 0.95. The $\overline{Z}$ of soot with C:H:O = 100:78:18 in atomic ratio is 4.17 and has a estimated density of 1.32(6) g/cc at 100 kPa and 300 K based on the adopted regression of $\rho_{0}$ = 0.317(15) $\overline{Z}$. This value is comparable to the proposed density of carbonaceous matter (1.4 g/cc) that coexists with hydrous silicates at 100 kPa and up to 900 K (neri2020).